OpenStudy (anonymous):

Help please! :) On a 50 question test, Mark got 60% of the 35 vocab questions correct and 80%of the 15 sentence structure questions correct. What percent of the 50 total questions did Mark get correct? 66% 61% 54% 18%

OpenStudy (anonymous):

as 60% of 35 vocab questions are correct => 60/100(35)=21 out of 35 vocab questions are correct. and 80% of 15 sentence structure are correct => 80/100(15)=12 out of 15 sentence structure are correct. totally 33 out of 50 questions are correct which means 33/50(100)= 66% of the questions are correct

OpenStudy (anonymous):

Thank you so much(: